OpenStudy (anonymous):

Diann bought a remnant of fabric that is 0.875 yards long. She is sewing an apron that takes 2/3 of a yard. How much fabric will Dian have left over after she sews the apron?

OpenStudy (anonymous):

7/8-2/3 = 21/24-16/24 = 5/24